Circular No. 8596

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V1047 CENTAURI = NOVA CENTAURI 2005 W. Liller, Viña del Mar, Chile, reports his discovery of an apparent

nova (mag ≈ 8.5) on two Technical Pan photographs taken on Sept. 1.031 UT with an 85-mm camera lens and an orange filter. The new object is located at α = 13^{h}20^{m}.8, δ = −62^{o}37′ (equinox 2000.0); nothing was seen at this location brighter than mag 11.0 on Aug. 12.050. CCD photometry on Sept. 4.028 yielded V = 8.83. C. Jacques and E. Pimentel, Belo Horizonte, Brazil, report the following precise position for this nova from a CCD image taken on Sept. 4.91 with a 0.3-m reflector that yielded V = 8.81: α = 13^{h}20^{m}49^{s}.74, δ = −62^{o}37′50′′.5. On Sept. 4.050, Liller obtained a red CCD spectrum (dispersion at Hα = 0.32 nm/pixel) with an objective grating and a 0.20-m Schmidt camera. The spectrum shows a moderately strong Hα emission line having a peak approximately twice the brightness of the neighboring continuum; its FWHM is 840 ± 50 km/s; no P-Cyg absorption is discernible. The close similarity of this spectrum with those of V5114 Sgr (N Sgr 2004) and V5116 Sgr (N Sgr 2005 No. 2) makes it virtually certain that the new object is indeed a classical nova. Visual magnitude estimates: Sept. 3.91, 9.7 (J. G. de S. Aguiar, Monte Mor, Brazil); 4.91, 8.9 (Aguiar); 4.927, 8.8 (R. Y. Shida, São Paulo, Brazil).

N. N. Samus, Institute of Astronomy, Russian Academy of Sciences, informs us that the designation V1047 Cen has been given to this nova.

2003 UB_{313}

J.-M. Petit, Observatoire de Besancon; M. Holman, Harvard-Smith-sonian Center for Astrophysics; and B. Gladman, University of British Columbia, report photometric observations of 2003 UB_{313} (cf. IAUC 8577). Holman obtained 120–240-s exposures during July 30.285–30.437 and 31.322–31.420 UT with the 6.5-m Baade telescope (+ Inamori Magellan Areal f/4.3 Camera and Spectrograph + Bessell R filter) at Las Campanas. Gladman obtained 30–120-s exposures during Aug. 2.278–2.435 and 3.330– 3.382 with the 8.2-m UT-1 Very Large Telescope (+ FORS2 camera + specialized Bessell R filter) at Cerro Paranal. Aperture photometry performed by Petit shows evidence of variability at the 0.015-mag level within two individual nights of observation. The night-to-night variations from July 30 to 31 and from Aug. 2 to 3 are of the same order. Data do not cover an arc long enough to determine a period, but if the variations are due to rotation, then the period exceeds 8 hr. However, the observed variation might be due to other effects, such as unknown color terms.

2005 September 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green